OpenStudy (chihiroasleaf):

(76 x 76 x 76 + 63 x 63 x 63) / (76 x 76 + 63 x 63 - 76 x 63) can we calculate the result without calculate each term ( I mean without finding 76 x 76 x 76 )

hartnn (hartnn):

yes.

OpenStudy (chihiroasleaf):

how?

hartnn (hartnn):

\(\large a^3+b^3 = (a+b)(a^2-ab+b^2)\)

hartnn (hartnn):

a+b= ?

hartnn (hartnn):

then put a= 76,b=63

hartnn (hartnn):

got that ^ ?

OpenStudy (chihiroasleaf):

yes..., got it..., :D thanks a lot...